A coffee merchant wants to make 25lbs of a blend of coffee to sell for $4.25 per pound. The blend is made by using a $6.00 grade and a $3.50 grade of coffee. How many pounds of each grade of coffee should be used?
Answer by checkley75(3666)

25*425=600X+350(25-X)
10625=600X+8750-350X
10625-8750=250X
1875=250X
X=1875/250
X=7.5 POUNDS OF $6.00 COFFEE.
25-7.5=17.5 POUNDS OF 43.50 COFFEE.
PROOF
10625=600*7.5+350*17.5
10625=4500+6125
10625=10625